OpenStudy (anonymous):

Two particles A and B have velocities 3i and vj respectively (in m/s). (a) Find rBjA; the position of B relative to A for all t given that rB/A(t = 0) = 10i + 5j (in metres).

OpenStudy (john_es):

One possible answer. Assuming they have constant velocities, then, \[r_B-r_A=(-3,v)t\] Add the initial condition, \[r_B-r_A=(10,5)+(-3,v)t\] And this should be the answer.